OpenStudy (anonymous):

Which sequence is generated by the function f(n + 1) = f(n) – 2 for f(1) = 10?

OpenStudy (hayleymeyer):

well, what do you think it is? first on all. :D

OpenStudy (anonymous):

no clue

OpenStudy (anonymous):

@Hayleymeyer u have to finish helping him lol

OpenStudy (hayleymeyer):

im not good at these D: @phi

OpenStudy (anonymous):

Which sequence is generated by the function f(n + 1) = f(n) – 2 for f(1) = 10? –10, –12, –14, –16, –18, ... –2, 8, 18, 28, 38, ... 8, 18, 28, 38, 48, ... 10, 8, 6, 4, 2, ...

OpenStudy (anonymous):

that's the whole question just don't know which answer it is

OpenStudy (anonymous):

The first term is f(1), that's 10. f(n + 1) = f(n) – 2 means you subtract 2 to get the next term

OpenStudy (phi):

f(n + 1) = f(n) – 2 for f(1) = 10 try to learn how to interpret this First, f(1) is short for "the first term" (that is what the 1 in f(1) means) so the first term in your answer should be a 10

OpenStudy (phi):

Can you learn that f(1) = some number that will be the first number in a list?

OpenStudy (anonymous):

thanks I got it